1/[x]-1/[3x]=2/9 where [x] is greatest  integer function find the value of x

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
if x is an integer then [x] = x and [3x] = 3x
 
hence 1/x – 1/3x = 2/9
 
2/3x = 2/9
3x = 9
 
hence x = 3